# Information-entropy-driven generation of material-agnostic datasets for machine-learning interatomic potentials

### v1

@@ Abstract
In contrast to their empirical counterparts, machine-learning interatomic potentials (MLIAPs) promise to deliver near-quantum accuracy over broad regions of configuration space. However, due to their generic functional forms and extreme flexibility, they can catastrophically fail to capture the properties of novel, out-of-sample configurations, making the quality of the training set a determining factor, especially when investigating materials under extreme conditions. In the present study, we propose a novel automated dataset generation method based on the maximization of the information entropy of the feature distribution, aiming at an extremely broad coverage of the configuration space in a way that is agnostic to the properties of specific target materials. The ability of the dataset to capture unique material properties is demonstrated on a range of unary materials, including elements with the fcc (Al), bcc (W), hcp (Be, Re and Os), and graphite (C) ground states. MLIAPs trained to this dataset are shown to be accurate over a broad range of application-relevant metrics, as well as extremely robust over very broad swaths of configurations space, even without dataset fine-tuning or hyper-parameter optimization, making the approach extremely attractive to rapidly and autonomously develop general-purpose MLIAPs suitable for simulations in extreme conditions.
@@ 1 Introduction
The high computational cost of ab initio electronic structure methods such as density functional theory (DFT) severely constrains the time and length-scales accessible to direct simulation. This has for decades motivated the development of computationally-efficient interatomic potentials that act as surrogates to direct quantum calculations. Early potentials relied on physically-motivated functional forms that involved a few empirical parameters that could be determined using only a small amount of training data (e.g., elastic constants, cohesive energy, lattice constants, etc.). Such simple functional forms rooted in physics typically resulted in robust potentials, at the cost of a generally significant compromise in accuracy [1]. In contrast, modern machine learning interatomic potentials (MLIAPs) instead invoke generic features (also known as descriptors) and/or flexible ML architectures to describe local atomic environments and capture their energetic contributions [2, 3, 4, 5, 6, 7, 8, 9, 10]. The high complexity and flexibility of these potentials dramatically increases the number of adjustable parameters compared to earlier generations of potentials, enabling unprecedented accuracy at the cost of a dramatic increase in the amount of required training data. Crucially, the lack of rigid physics-based underpinnings potentially leads to poor extrapolative capabilities, making MLIAPs susceptible to catastrophic failures in regimes where training data is either scarce or missing [11, 12], even when test errors on hold-out data suggests that the potentials should be very accurate [12]. Observations suggest that assembling diverse datasets with high energy structures that are far from equilibrium significantly improves the stability of simulations [12]. A similar conclusion was reached in Ref. [13], where serious limitations in the robustness of models trained to hand-curated and narrowly tailored datasets were observed. This highlights the critical importance of training MLIAPs to very diverse data, especially in applications to extreme conditions where it is very difficult to a priori predict the regions of configuration space that will be visited by simulations. This recognition has led to the development of a number of approaches aiming at dramatically improving dataset curation compared to the manual domain-expertise-driven approaches that were previously common.
@@
Perhaps the most well-known approach in this class is active learning, where iteratively-refined potentials are used to drive further exploration of the configuration space, repeating the training/exploration cycle multiple times. These techniques rely on a combination of diverse initialization and/or efficient sampling algorithms (e.g., molecular dynamics [14], evolutionary structural searches [15], local optimization [16, 17], Bayesian regression [18], dimensionality reduction [19]) in order to ensure broad coverage of configuration space. Methods in this class often also rely on uncertainty quantification (e.g., using extrapolative score, d-optimality [20, 21, 22, 23], Bayesian regression [17, 24, 25]) to flag configurations where the current potential is deemed unreliable for inclusion in the next iteration of the training set. These approaches can be an extremely powerful way to enrich datasets with targeted information directly relevant to certain conditions or applications. However, given the rugged nature of actual energy surfaces, the diversity of the dataset produced by active learning can be expected to be significantly affected by the details of the initialization and sampling procedures. Care must therefore be taken if one aims at a very broad exploration as energetically or entropically hard-to-reach regions can be expected to be the norm rather than the exception.
@@
In the following, we generalize the aforementioned entropy optimization approach to generate a compact, yet ultra-diverse, material-agnostic dataset that can be applied to train robust unary MLIAPs across the periodic table. The power of this approach is demonstrated by training potentials in the family of Atomic Cluster Expansion (ACE) [10] to a diverse set of elements (Be, C, Al, W, Re, Os), thoroughly validating the results using physical properties that are not explicitly targeted in the training data. The proposed approach provides a systematic, simple, and fast approach to MLIAP development that alleviates the need for fine-tuned material-specific datasets, while retaining the type of exceptional transferability and predictive capabilities that is required to investigate materials in broad range of conditions. Although not pursued here, this method can easily be combined with traditional active learning techniques to further fine-tune the models in applications-specific regions of configuration space.
The manuscript is organized as follows. In Sec. 2, we explain the workflow of the entropy maximization approach followed by a comparison of the diversity of the resulting dataset to those from literature in Sec. 2.2. In Sec. 2.4, we extensively validate the performance of a W ACE potential fitted to Rev-EM and analyze the accuracy/robustness tradeoff of W ACE models fitted to the Rev-EM dataset in Sec. 2.5. The material agnostic nature of the Rev-EM dataset is demonstrated in Sec. 2.6 by validating ACE models for Be, C, Al, Re, and Os. Details of the computational procedure are given in Sec. 5.
@@ 2.1 Entropy maximization approach
which, up to irrelevant additive and multiplicative constants, is simply proportional to the logarithm of the determinant of the feature covariance matrix; in an abuse of language, we will refer to $\log\det(\Sigma)$ as the information entropy in the following. This quantity has a simple interpretation in terms of the sum of the logarithm of the variance of the feature distribution along all principal directions: high feature entropies hence corresponds to broad distributions in feature space. Note that within this approximation, maximizing the dataset entropy is akin to a d-optimality experiment design procedure, which is known to optimize the differential Shannon information content of the parameter estimates. D-optimality was introduced in the context of active learning of MLIAPs by Podryabinkin and Shapeev [20], where the determinant of the feature covariance matrix was used as a scoring function for candidate configurations. The original "passive" approach is here generalized to an "active" variant where atomic configurations are explicitly generated so as to maximize this objective function.
In practice, $\Sigma$ is given by the empirical covariance of the features over a given dataset. In the following, we consider two flavors, which are referred to as per-atom and per-configuration respectively. In the per-atom case, each atom $i$ contributes a feature vector $B_{i}$, while in the per-configuration variant, each configuration $\alpha$ contributes a feature vector $\hat{B}_{\alpha}$ which is the arithmetic mean of the atomic $B_{i}$ within the configuration. This distinction is made by analogy to the training of an hypothetical MLIAP which would be linear in the $B_{i}$. In this case, the predicted energy per atom would be proportional to the mean of the $B_{i}$ over all atoms in a given configuration. We empirically observe that maximizing the entropy of per-atom features tends to generate internally diverse configurations, while maximizing the entropy of per-configuration features tends to, perhaps surprisingly, generate internally ordered configurations. This can be understood qualitatively in terms of self-averaging where, by definition, the mean of per-atom features must lie within the envelope of the individual per-atom features. Generating internally diverse configurations is therefore in tension with the generation of extreme values of the mean, which is comparatively easier to achieve when all per-atom features take on similar values, hence resulting in internally ordered configurations.
@@ 2.2 Dataset diversity
In order to quantitatively compare the diversity of Rev-EM with other curation approaches, the bispectrum component feature entropy of a number of publicly available datasets was estimated after volumetrically rescaling all configurations to a common reference density. These datasets include Poul-Mg [26], Bernstein-Ti [16], Smith-Al [14], Lysogorskiy-Cu [31], Song-W [32], Owen-W [33] and Karabin-W, a subset of 20,000 configuration from Ref. [27]. See Sec. 2.1 for more details on the different datasets. Given that the bispectrum components exhibit a strong volume dependence, only configurations within $\pm$25of the equilibrium density were considered to avoid the specific range of densities in each set dominating the comparison. Note that this does not fully correct for this effect, as the distribution within each dataset will still differ. The results are reported in Fig. 3. A comparison of the original un-truncated datasets is presented in Fig. S2, which shows that the extremely high density configurations included in Poul-Mg confer it a very high information entropy. In addition to recognizing the caveats above, we emphasize that the information entropy should not be construed as an absolute measure of the quality of the datasets and of the MLIAPs derived from them, as each set was created with different applications, purposes, and constraints in mind. Instead, it should be simply taken as a reflection of the diversity of the configurations included in each dataset, which can be beneficial but also potentially detrimental, depending on the application.
@@ 2.3 Training MLIAPs to Rev-EM
We now demonstrate that MLIAPs trained to Rev-EM nonetheless perform extremely well on a broad range of metrics of interest to computational materials scientists. This demonstration proceeds in three steps. First, in Sec. 2.4, we characterize the performance of a tungsten ACE potential fitted to Rev-EM using an extensive set of common benchmarks. Second, Sec. 2.5 analyzes the accuracy/robustness tradeoff resulting from this dataset in terms of accuracy on low-energy structures vs transferability to vast regions of configuration space. Third, Sec. 2.6 demonstrates the transferability of Rev-EM across materials, including Al, C, Be, Re, and Os.
@@
Rev-EM was randomly partitioned into around 8000 training configurations and around 22,000 test configurations (the exact number depending on the energy/force cutoffs for each element). More details on the training and test configurations for all elements can be found in Secs. 5.4 and 5.5. Such a large testing set was purposefully chosen so as to provide a very challenging assessment of the robustness of the potentials on large numbers of ultra-diverse configurations. The reported test errors are therefore expected to constitute a good measure of the transferability of the MLIAPs to genuinely "unseen" and "surprising" configurations, which in general cannot be expected with random hold-out data from low-diversity datasets, where strong correlations between train and test sets would be expected. When training production models, it is advisable to increase the fraction of training to more typical values.
@@ 2.4 A case study for W
While the reported errors are relatively large compared to fine-tuned MLIAPs, it is important to note that a reweighting scheme was applied during training to give more importance to low energy and force structures that are usually prioritized in applications (see Sec. 5.5); this is reflected by errors being generally lower at low energies and forces. For example, RMSE errors are significantly lower at 23.83 meV/atom and 216.25 meV/Å  for energies and forces, respectively, for configurations within 1 eV of the lowest-energy structure. Furthermore, transition metals are known to be particularly difficult to describe with MLIAPs, due to the complexity of their electronic structure around the Fermi level [33]. As shown in Table 1, the test errors on of ACE-2 on DFT datasets for W available from the literature show significantly lower errors than on holdout data from Rev-EM, supporting the idea that these are functionally low-energy subsets of Rev-EM. Note that the original energies and forces reported by the respective studies were used to evaluate the testing errors; different DFT settings could therefore lead to a modest overestimation of ACE-2 errors.
@@
Fig. 6 shows that ACE-2 accurately predicts the relative formation energies, formation volumes and bulk moduli of a set of 18 crystal structures ranging from the known BCC ground state to high-energy structures such as A4 (which is about 2.5 eV/atom higher in energy). None of these crystal structures were manually introduced into Rev-EM and that some of them (including TCP and laves phases) have unit cells that are larger than the largest configuration in Rev-EM (25 atoms). The largest relative errors are observed for the bulk modulus, with errors less than about 5% on all phases. Agreement with cubic elastic constants is similarly good, as shown in Tab. 2, although an error of about 23% is observed for $C_{12}$. Transformation paths (generated following the procedure presented in Refs. [37, 38, 39]) between some of these same crystal structures are also extremely well captured, as shown in Fig. 7, demonstrating that ACE-2 accurately captures the energy landscape between the local minima corresponding to each structure.

### v2

@@ Abstract
In contrast to their empirical counterparts, machine-learning interatomic potentials (MLIAPs) promise to deliver near-quantum accuracy over broad regions of configuration space. However, due to their generic functional forms and extreme flexibility, they can catastrophically fail to capture the properties of novel, out-of-sample configurations, making the quality of the training set a determining factor, especially when investigating materials under extreme conditions. We propose a novel automated dataset generation method based on the maximization of the information entropy of the feature distribution, aiming at an extremely broad coverage of the configuration space in a way that is agnostic to the properties of specific target materials. The ability of the dataset to capture unique material properties is demonstrated on a range of unary materials, including elements with the fcc (Al), bcc (W), hcp (Be, Re and Os), graphite (C), and trigonal (Sb, Te) ground states. MLIAPs trained to this dataset are shown to be accurate over a range of application-relevant metrics, as well as extremely robust over very broad swaths of configurations space, even without dataset fine-tuning or hyper-parameter optimization, making the approach extremely attractive to rapidly and autonomously develop general-purpose MLIAPs suitable for simulations in extreme conditions.
@@ 1 Introduction
The high computational cost of ab initio electronic structure methods such as density functional theory (DFT) severely constrains the time and length-scales accessible to direct simulations. This has for decades motivated the development of computationally-efficient interatomic potentials that act as surrogates to quantum calculations. Early potentials relied on physically-motivated functional forms that involved a few empirical parameters that could be determined using only a small amount of training data (e.g., elastic constants, cohesive energy, lattice constants, etc.). Such simple functional forms rooted in physics typically resulted in robust potentials, at the cost of a generally significant compromise in accuracy [1]. In contrast, modern machine learning interatomic potentials (MLIAPs) instead invoke generic features (also known as descriptors) and/or flexible ML architectures to describe local atomic environments and capture their energetic contributions [2, 3, 4, 5, 6, 7, 8, 9, 10]. The high complexity and flexibility of these potentials dramatically increases the number of adjustable parameters compared to earlier generations of potentials, enabling unprecedented accuracy at the cost of a very significant increase in the amount of required training data. Crucially, the lack of rigid physics-based underpinnings potentially can lead to poor extrapolative capabilities, making MLIAPs susceptible to catastrophic failures in regimes where training data is either scarce or missing [11, 12], even when test errors on hold-out data suggests that the potentials should be very accurate [12]. Observations suggest that assembling diverse datasets with high energy structures that are far from equilibrium significantly improves the stability of simulations [12]. A similar conclusion was reached in Ref. [13], where serious limitations in the robustness of models trained to hand-curated and narrowly tailored datasets were observed. This highlights the critical importance of training MLIAPs to very diverse data, especially in applications to extreme conditions where it is very difficult to a priori predict the regions of configuration space that will be visited by simulations. This recognition has led to the development of a number of approaches aiming at dramatically improving dataset curation compared to the manual domain-expertise-driven approaches that were previously common.
(a) DFT
(b) ACE [14]
(c) ACE (this work)
(d) SNAP [15]
(e) ANI-1xnr [16]
(f) GAP [17]
Figure 1: Transferability of different MLIAPs for C to nine crystal structures.
@@
Perhaps the most well-known approach in this class is active learning, where iteratively-refined potentials are used to drive further exploration of the configuration space, repeating the training/exploration cycle multiple times. These techniques rely on a combination of diverse initialization and/or efficient sampling algorithms (e.g., molecular dynamics [18, 19, 20], metadynamics [21], uncertainty-biased MD [22, 23], evolutionary structural searches [24, 25], local structure optimization and diversity based selection [25], clustering [26], dimensionality reduction [27], etc.) to ensure broad coverage of configuration space. Methods in this class often also rely on uncertainty quantification (e.g., using extrapolative score, d-optimality [28, 29, 30, 31, 24], Bayesian regression [20, 22, 32, 33], etc.) to flag configurations where the current potential is deemed unreliable for inclusion in the next iteration of the training set. These approaches can be a powerful way to enrich datasets with targeted information directly relevant to certain conditions or applications. However, given the rugged nature of actual energy surfaces, the diversity of the dataset produced by active learning can be expected to be significantly affected by the details of the initialization and sampling procedures. Care must therefore be taken if one aims at a very broad exploration as energetically or entropically hard-to-reach regions can be expected to be the norm rather than the exception.
@@
To exemplify these transferability challenges, we tested several popular MLIAPs developed for C using a range of different data generation strategies and base ML architectures. Energy-volume curves for different crystal structures are reported in Fig. 1. We emphasize that the developers of these models (Spectral Neighbor Analysis Potential (SNAP) [15], ANI-1xnr [16] and Gaussian Approximation Potential (GAP) [17]) in general made no claims on accuracy of the models for these phases, except in the case of the Atomic Cluster Expansion (ACE) model from Ref. [14] (c.f., Fig. 1(b)) where such data was actually used in training. The SNAP, ANI-1xnr and GAP models show numerous spurious features in their energy landscapes. This indicates that even with expressive ML architectures and careful data curation by domain experts, ML potentials cannot be expected to capture properties that they were not explicitly trained to and are hence prone to unphysical behaviors away from regions covered by training data. In contrast, as will be shown below, the potentials trained in this work (c.f., Fig. 1(c)) captures the behavior of a broad range of phases that were not explicitly injected in the training set. This indicates that one-shot diverse-by-design data generation approaches can in fact efficiently constrain the behavior of flexible ML potentials over broad swaths of the feature space without requiring external guidance from domain experts or iterative improvement with active learning, opening the door to the straightforward and principled generation of training sets that confer extreme robustness to ML potentials.
In the following, we generalize the aforementioned entropy optimization approach to generate a compact, yet ultra-diverse, material-agnostic dataset that can be applied to train robust unary MLIAPs across the periodic table. The power of this approach is demonstrated by training potentials in the family of ACE [10] to a diverse set of elements (Be, C, Al, Sb, Te, W, Re, Os), thoroughly validating the results using physical properties that are not explicitly targeted in the training data. The proposed approach provides a systematic, simple, and fast approach to MLIAP development that alleviates the need for fine-tuned material-specific datasets, while retaining the type of exceptional transferability and predictive capabilities that is required to investigate materials in broad range of conditions. Although not pursued here, this method can easily be combined with traditional active learning techniques to further fine-tune the models in applications-specific regions of configuration space.
The manuscript is organized as follows. In Sec. 2, we explain the workflow of the entropy maximization approach followed by a comparison of the diversity of the resulting dataset to those from literature in Sec. 2.2. In Sec. 2.4, we extensively validate the performance of an ACE potential for W that is fitted to the newly generated dataset and analyze the accuracy/robustness tradeoff of the resulting W ACE models in Sec. 2.5. The material agnostic nature of the current dataset is demonstrated in Sec. 2.6 by validating ACE models for Be, C, Al, Sb, Te, Re, and Os. The exceptional robustness of the ACE models for Al, W, and C is further corroborated by cross-validating on multiple datasets available from the literature in Sec. 2.7. Details of the computational procedure are given in Sec. 5.
@@ 2.1 Entropy maximization approach
which, up to irrelevant additive and multiplicative constants, is simply proportional to the logarithm of the determinant of the feature covariance matrix; in an abuse of language, we will refer to $\log\det(\Sigma)$ as the information entropy in the following. This quantity has a simple interpretation in terms of the sum of the logarithm of the variance of the feature distribution along all principal directions: high feature entropies hence corresponds to broad distributions in feature space. Note that within this approximation, maximizing the dataset entropy is akin to a d-optimality experiment design procedure, which is known to optimize the differential Shannon information content of the parameter estimates. This correspondence with d-optimality suggests that whether or not the normal approximation is an accurate representation of the actual feature distribution, maximizing Eq. 3 will nonetheless maximize a well-defined measure of dataset diversity. This is an important point given that, as will be shown below in Fig. 5, certain feature distributions show clear signs of non-normality. D-optimality was introduced in the context of active learning of MLIAPs by Podryabinkin and Shapeev [28], where the determinant of the feature covariance matrix was used as a scoring function for candidate configurations. The original "passive" approach is here generalized to an "active" variant where atomic configurations are explicitly generated so as to maximize this objective function.
In practice, $\Sigma$ is given by the empirical covariance of the features over a given dataset. In the following, we consider two flavors, which are referred to as per-atom and per-configuration respectively. In the per-atom case, each atom $i$ contributes a feature vector $B_{i}$, while in the per-configuration variant, each configuration $\alpha$ contributes a feature vector $\hat{B}_{\alpha}$ which is the arithmetic mean of the atomic $B_{i}$ within the configuration. This distinction is made by analogy to the training of an hypothetical MLIAP which would be linear in the $B_{i}$. In this case, the predicted energy per atom would be proportional to the mean of the $B_{i}$ over all atoms in a given configuration. We empirically observe that maximizing the entropy of per-atom features tends to generate internally diverse configurations, while maximizing the entropy of per-configuration features tends to, perhaps surprisingly, generate internally ordered configurations. This can be understood qualitatively in terms of self-averaging where, by definition, the mean of per-atom features must lie within the envelope of the individual per-atom features. Generating internally diverse configurations is therefore in tension with the generation of extreme values of the mean, which is comparatively easier to achieve when all per-atom features take on similar values, hence resulting in internally ordered configurations. We however note that any other, potentially non-linear, combination of atomic features could in principle be used as a basis for the entropy optimization procedure.
@@ 2.2 Dataset diversity
In order to quantitatively compare the diversity of Rev-EM with other curation approaches, the bispectrum component feature entropy of a number of publicly available datasets was estimated after volumetrically rescaling all configurations to a common (average) reference density. These datasets include Poul-Mg [34], Bernstein-Ti [25], Smith-Al [18], Lysogorskiy-Cu [39], Song-W [40], Owen-W [41] and Karabin-W, a subset of 20,000 configuration from Ref. [35]. See Sec. 2.1 for more details on the different datasets. Given that the bispectrum components exhibit a strong volume dependence, only configurations within $\pm$25of the equilibrium density were considered to avoid the specific range of densities in each set dominating the comparison. Note that this does not fully correct for this effect, as the distribution of volumes within each dataset will still differ. The results are reported in Fig. 4. A comparison of the original un-truncated datasets is presented in Fig. S2, which shows that the extremely high density configurations included in Poul-Mg confer it a very high information entropy. In addition to recognizing the caveats above, we emphasize that the information entropy should not be construed as an absolute measure of the quality of the datasets and of the MLIAPs derived from them, as each set was created with different applications, purposes, and constraints in mind. Instead, it should be simply taken as a reflection of the diversity of the configurations included in each dataset, which can be beneficial but also potentially detrimental, depending on the application.
@@ 2.3 Training MLIAPs to Rev-EM
We now demonstrate that MLIAPs trained to Rev-EM nonetheless perform extremely well on a broad range of metrics of interest to computational materials scientists. This demonstration proceeds in three steps. First, in Sec. 2.4, we characterize the performance of a tungsten ACE potential fitted to Rev-EM using an extensive set of common benchmarks. Second, Sec. 2.5 analyzes the accuracy/robustness tradeoff resulting from this dataset in terms of accuracy on low-energy structures vs transferability to vast regions of configuration space. Third, Sec. 2.6 demonstrates the transferability of Rev-EM across materials, including Al, C, Be, Re, Os, Te, and Sb.
@@
Rev-EM was randomly partitioned into around 8000 training configurations and around 22,000 test configurations (the exact number depending on the energy/force cutoffs for each element). More details on the training and test configurations for all elements can be found in Secs. 5.4 and 5.5. Such a large testing set was purposefully chosen so as to provide a very challenging assessment of the robustness of the potentials on large numbers of ultra-diverse configurations. The reported test errors are therefore expected to constitute a good measure of the transferability of the MLIAPs to genuinely "unseen" and "surprising" configurations, which in general cannot be expected with random hold-out data from low-diversity datasets, where strong correlations between train and test sets would be expected. When training production models, it is advisable to increase the fraction of training to more typical values, which can be expected to further improve accuracy.
@@ 2.4 A case study for W
While the reported errors are relatively large compared to fine-tuned MLIAPs, it is important to note that a reweighting scheme was applied during training to give more importance to low energy and force structures that are usually prioritized in applications (see Sec. 5.5); this is reflected by errors being generally lower at low energies and forces. For example, RMSE errors are significantly lower at 23.83 meV/atom and 216.25 meV/Å  for energies and forces, respectively, for configurations within 1 eV of the lowest-energy structure. Furthermore, transition metals are known to be particularly difficult to describe with MLIAPs, due to the complexity of their electronic structure around the Fermi level [41]. It finally should be recognized that reported test errors are intimately related to the diversity of the test sets and to the extent to which training and testing sets are correlated to each other. As such, it is generally impossible to rigorously compare the accuracy of different potentials using test errors measured on different sets. Indeed, as shown in Table 1, the test errors of ACE-2 on other DFT datasets for W available from the literature show significantly lower errors than on holdout data from Rev-EM, supporting the idea that these are functionally low-energy subsets of Rev-EM. As will be shown below, the reverse is typically observed when testing potentials trained to other datasets to Rev-EM, i.e., they tend to exhibit significantly higher errors than reported on their original test set. Note that the original energies and forces reported by the respective studies were used to evaluate the testing errors; different DFT settings could therefore lead to a overestimation of the errors when mixing training and test sets from different studies.
@@
Fig. 7 shows that ACE-2 accurately predicts the relative formation energies, formation volumes and bulk moduli of a set of 18 crystal structures ranging from the known BCC ground state to high-energy structures such as A4 (which is about 2.5 eV/atom higher in energy). None of these crystal structures were manually introduced into Rev-EM and that some of them (including TCP and laves phases) have unit cells that are larger than the largest configuration present in Rev-EM (25 atoms). The largest relative errors are observed for the bulk modulus, with errors less than about 5% on all phases. Agreement with cubic elastic constants is similarly good, as shown in Tab. 2, although an error of about 23% is observed for $C_{12}$. Transformation paths (generated following the procedure presented in Refs. [45, 46, 47]) between some of these same crystal structures are also extremely well captured, as shown in Fig. 8, demonstrating that ACE-2 accurately captures the energy landscape between the local minima corresponding to each structure.